OpenStudy (anonymous):

For the given quadratic function use the method of "completing the square" to convert to a-h-k form, and then identify the vertex. r(s)= -7s^2 +42s-33=_______________________ has its vertex at (______,________) .

OpenStudy (precal):

factor out the -7in the first two terms, then complete the square, don't forget to add and subtract it and don't forget the -7

OpenStudy (anonymous):

there is an easier way than completing the square you know it is going to look like \[r(s)=-7(r-h)^2+k\] and the vertex will be \((h,k)\) vertex is always at \(-\frac{b}{2a}\) which in your case is \(-\frac{42}{2\times (-7)}=3\)

OpenStudy (anonymous):

so you know it is going to be \[r(s)=-7(s-3)^2+k\] and to find \(k\) replace \(s\) by \(3\) in the original expression and compute \(r(3)\)